# A classification of Fourier summation formulas and crystalline measures

### v1

@@ Abstract.
We completely classify Fourier summation formulas, and in particular all crystalline measures, under quadratic decay. Our classification employs techniques from almost periodic functions, Hermite-Biehler functions, de Branges spaces and Poisson representation. We show how our classification generalizes recent results of Kurasov & Sarnak and Olevskii & Ulanovskii. As an application, we give a new classification result for nonnegative crystalline measures that are bounded away from zero on their support. In the Appendix we attempt to give a systematic account of the several types of Fourier summation formulas, including a new construction using eta-products, generalizing an old example of Guinand.
@@ 1. Introduction
To understand the atomic structure of a crystal (of some material) one can fire a beam of electrons into the object and observe the electron diffraction pattern on a screen [^1] This technique is called high resolution electron microscopy (HRTEM), but another common technique is X-ray microscopy.. In most materials one observes a periodic diffraction pattern, such as a two-dimensional lattice with $6$-fold symmetry, for instance as in Tantalum pentoxide. This is not the case with Holmium–magnesium–zinc, which one can observe a $10$-fold symmetry, hence a non-periodic diffraction pattern. The first to observe this experimentally was materials scientist Dan Shechtman, in 1982, and today these materials are called quasicrystals. He was awarded the Nobel Prize in Chemistry in 2011 for his breakthrough.
@@
Mathematically, if $\mu$ represents the crystal (and so $\mu\geq 0$) then what we see in the diffraction pattern is effectively $|\widehat{\mu}|^{2}$. In this way, after Shechtman’s paper, many mathematicians started wondering about what different types of measures $\mu$ with locally finite support and spectrum are there, and what kind of geometrical properties of the spectrum are allowed under various conditions on $\mu$? The obvious big question being:
Can on one classify all crystalline measures?
For now on, we say a measure $\mu$ is crystalline if it has locally finite support and its Fourier transform $\widehat{\mu}$ also has locally finite support [^2] Such measures are sometimes also called quasicrystals, or Fourier quasicrystals, depending on the author, context and decay conditions. We prefer to stay away from the physics nomenclature.. Freeman Dyson addressed this question in his famous paper [14], stating that the classification of crystalline measures is a very difficult mathematical problem, but made the “outrageous suggestion” it could potentially give a proof of the Riemann Hypothesis [^3] We do not share the same belief.. The notorious paper of Guinand [18], which inspired Dyson, grasps such classification, although mostly overlooked. Nowadays there is unified framework to construct measures $\mu$ whose spectrum fits the data observed in the diffraction patterns of physical quasicrystals, these are called model sets (or Meyer sets) [30, 31, 15]. However, the mathematical question still is wide open.
@@
In this manuscript we deal only with the $1$-dimensional case. In this scenario, the first recent breakthrough occurred in the work of Lev & Olevskii [24, 25], where they classify classical Poisson summation as the only measures that have uniformly discrete support and spectrum. Guinand’s example [18, p. 265] shows the theorem is tight (we generalize his construction in the Appendix). Recently, the field is very active, and the amazing works of Radchenko & Viazovska [33], Kurasov & Sarnak [21], Olevskii & Ulanovskii [32] and Alon et. al. [2, 1], but also, the not so recent works of Bohr [12], Besicovitch [7], Guinand [18], de Branges [9] and Meyer [29], deeply inspired this paper.
@@
In this paper we completely classify (what we call) Fourier summation formulas, and in particular we classify crystalline measures that have quadratic decay. As far as we know, this is the first time such classification results are proven. Our classification involves heavily the theory of almost periodic analytic functions [7] and de Branges spaces of entire functions [9]. A Fourier summation formula should be one where the identity
@@
holds, at the bear minimum, for all functions $\varphi\in C^{\infty}_{c}(\mathbb{R})$, where $\{\lambda_{n}\}_{n\in\mathbb{N}}\subset\mathbb{R}$ is some sequence of real nodes, $\{a(\lambda_{n})\}_{n\in\mathbb{N}}\subset\mathbb{C}$ are some complex coefficients and $\mu$ is some complex-valued measure. There is a zoo of formulas of such type which we describe in the Appendix, from ancient and modern times. We choose the above form because: (1) We want a summation formula, so there must be a sum somewhere in the definition and such summation must be singular with respect to Lebesgue measure; (2) Allowing the right hand side above to have a vanishing singular part is boring, as otherwise any $L^{1}$-function would give rise to a Fourier summation formula.
@@ 1.1. Overview
Theorems 1 and 2 are our main general classification results and we later specialize them for crystalline measures in Theorems 3 and 4. Finally, we derive a new characterization Theorem 5 for crystalline measures $\mu$ such that $\mu|_{\operatorname{\mathrm{supp}}(\mu)}\geq\varepsilon$.
@@ 2. Main Results
We will have to introduce several things before we can state our first main result [^4] So bear with me. We say that a complex-valued measure $\mu$ on $\mathbb{R}$ is locally finite if its total variation
@@
is finite on every compact set $K\subset\mathbb{R}$, or equivalently, if there are nonnegative locally finite Borel measures $\mu_{j}$ such that $\mu=\mu_{1}-\mu_{2}+i(\mu_{3}-\mu_{4})$. We say that $\mu$ is tempered if it is a complex-valued locally finite measure and there exists a tempered distribution $u\in\mathcal{S}^{\prime}(\mathbb{R})$ such that $u(\varphi)=\int_{\mathbb{R}}\varphi\mathrm{d}\mu$ for all $\varphi\in C^{\infty}_{c}(\mathbb{R})$. Some authors might not concord with this definition, but we are following the setup of [4] (and the references therein). We say that $\mu$ is strongly tempered if $\varphi\in L^{1}(|\mu|)$ for all $\varphi\in\mathcal{S}(\mathbb{R})$. By [4, Prop. 2.5], this is equivalent to the degree of $\mu$
@@
being <absent<$\infty$< ∞. In particular, the functional $\varphi\mapsto\int_{\mathbb{R}}\varphi\mathrm{d}\mu$ defines a tempered distribution, hence $\mu$ is tempered. The implication [tempered] $\Rightarrow$ [strongly tempered] is not true. Indeed, the measure $\mu=\sum_{n\geq 0}3^{n^{2}}({\boldsymbol{\delta}}_{n}-{\boldsymbol{\delta}}_{n+4^{-n^{2}}})$ is not strongly tempered since $\int 2^{-t^{2}}\mathrm{d}|\mu|(t)\geq\sum_{n\geq 0}3^{n^{2}}2^{-n^{2}}=\infty$, but it is tempered [^5] Whenever it is convenient, we write $A\ll B$ to mean that there is $C>0$ such that $|A|\leq C|B|$. $|\sum_{n\geq 0}3^{n^{2}}(\varphi({n})-\varphi({n+4^{-n^{2}}}))|\ll\|\varphi^{\prime}\|_{\infty}$.
We say that a function $a:\mathbb{R}\to\mathbb{C}$ is locally summable if the support set $\operatorname{\mathrm{supp}}(a):=\{\lambda\in\mathbb{R}:a(\lambda)\neq 0\}$ is at most countable and for some enumeration $\operatorname{\mathrm{supp}}(a)=\{\lambda_{n}\}_{n\geq 1}$ we have
@@
We say $(\mu,a)$ is a Fourier summation pair $(\mathrm{FS}$-pair)normal-))) if $\mu$ is a strongly tempered measure, $a:\mathbb{R}\to\mathbb{C}$ is locally summable and
@@
for $j=1,2$. Since for any $\varphi\in C^{\infty}_{c}(\mathbb{R})$ we have $\varphi=\varphi_{1}-i\varphi_{2}$ where each $\varphi_{j}$ is antipodal, and each $\varphi_{j}={\rm Re}\,\varphi_{j}-i{\rm Im}\,\varphi_{j}$, where ${\rm Re}\,\varphi_{j}$ is real and even and $i{\rm Im}\,\varphi_{j}$ is imaginary and odd (hence both are antipodal as well), we deduce by linearity that $(\mu_{j},a_{j})$ is a real-antipodal $\mathrm{FS}$-pair for $j=1,2$. As an example, note that $\sum_{n\in\mathbb{Z}}i^{n}\varphi(n)=\sum_{n\in\mathbb{Z}}\widehat{\varphi}(n-1/4)$ has the real-antipodal splitting $\sum_{n\in\mathbb{Z}}(-1)^{n}\varphi(2n)=\frac{1}{2}\sum_{n\in\mathbb{Z}}\left(\widehat{\varphi}(n-1/4)+\widehat{\varphi}(n+1/4)\right)$ and $\sum_{n\in\mathbb{Z}}(-1)^{n}\varphi(2n-1)=\frac{1}{2i}\sum_{n\in\mathbb{Z}}\left(\widehat{\varphi}(n+1/4)-\widehat{\varphi}(n-1/4)\right)$.
@@
for some $c>0$. We let +=$\mathbb{C}^{+}=\{z=x+iy:y>0\}$. Following Besicovitch [^6] Besicovitch considers the half-plane ${\rm Re}\,z>0$ instead of ${\rm Im}\,z>0$, so one has to mentally perform a rotation in order to use his results. [7, Ch. 3], we say a holomorphic function $f:\mathbb{C}^{+}\to\mathbb{C}$ is almost periodic if for every $0<\alpha<\beta<\infty$ and $\varepsilon>0$ there is a relatively dense [^7] A set $\tau\subset\mathbb{R}$ is relatively dense if there exists $l>0$ such that $(x,x+l)\cap\tau\neq\emptyset$ for every $x\in\mathbb{R}$. set of $\varepsilon$-translations $\tau\subset\mathbb{R}$ such that
@@
${\rm spec}(f):=\{\lambda\in\mathbb{R}:\mathbb{E}f(\lambda)\neq 0\}.$
We say a holomorphic function $g:\mathbb{C}^{+}\to\mathbb{C}$ is of bounded type [^8] This is the same to say that $\log|g(z)|$ has a nonnegative harmonic majorant. if there are two bounded holomorphic functions $P,Q:\mathbb{C}^{+}\to\mathbb{C}$ such that $g=P/Q$.
The following is the first main result of this paper.
@@
    $\exp(f)$ is of bounded type and $\limsup_{y\to\infty}\frac{{\rm Re}\,f(iy)}{y}=0$.
@@
Note that condition ($*{}^{*}$) is equivalent to $\mathbb{E}f(\lambda)$ being locally summable.
@@
We claim that if $(\mu,a)$ is an $\mathrm{FS}$-pair such such that $\mu\geq 0$ and $a(\cdot)$ has exponential growth then $\deg(\mu)\leq 2$. Indeed, let $\varphi\in C^{\infty}_{c}(\mathbb{R})$ be even, $0\leq\varphi\leq 1$, $\int\varphi=\varphi(0)=1$ and $\operatorname{\mathrm{supp}}(\varphi)\subset(-1,1)$. Let $\varphi_{\varepsilon}(x)=\varphi(x/\varepsilon)/\varepsilon$ and $\psi_{\varepsilon}(x)=\widehat{\varphi}(x/\varepsilon)/\varepsilon$ for $10<\varepsilon<$ 1. Let $g(x)=e^{-2\pi C|x|}$, with $C>0$ sufficiently large, and note $\widehat{g}(\xi)=\frac{C}{\pi(C^{2}+\xi^{2})}$. We have
$\int_{\mathbb{R}}(\widehat{g}\widehat{\varphi}_{\varepsilon})*\psi_{\varepsilon}(t)\mathrm{d}\mu(t)=\sum_{\lambda\in\mathbb{R}}a(\lambda)g*\varphi_{\varepsilon}(\lambda)\widehat{\psi}_{\varepsilon}(\lambda).$
It is not hard to show that $|g*\varphi_{\varepsilon}(\lambda)\widehat{\psi}_{\varepsilon}(\lambda)|\leq e^{-2\pi C(|\lambda|-1)}$. Thus, we conclude the right hand side above is uniformly bounded for $10<\varepsilon<$ 1. Since $\mu\geq 0$ and $(\widehat{g}\widehat{\varphi}_{\varepsilon})*\psi_{\varepsilon}\to\widehat{g}$ pointwise, as $\varepsilon\to 0$, we can apply Fatou’s Lemma to deduce $\int\frac{\mathrm{d}\mu(t)}{C^{2}+t^{2}}<\infty$.
Remark 3.
Let $(\mu,a)$ be an $\mathrm{FS}$-pair such that $\mu\geq 0$ and $a(\cdot)$ satisfies
$\sum_{\lambda\in\mathbb{R}}|a(\lambda)|e^{-|\lambda|/\sigma(\lambda)}<\infty,$ (5)
for some function $\sigma:\mathbb{R}\to[1,\infty)$ such that $\int_{|x|\geq 1}\frac{1}{x\sigma(x)}\mathrm{d}x<\infty$. We claim there is $C>0$ such that $\mu([x,x+1])\leq C$ for all $x\in\mathbb{R}$, hence $\deg(\mu)\leq 1$. For instance, one could take $\sigma(\lambda)=\log(10+|\lambda|)(\log\log(10+|\lambda|))^{2}$. Indeed, we only need to consider Ingham’s construction [19]. This is an auxiliary even function $\varphi\in\mathcal{S}(\mathbb{R})$ such that $\varphi\geq 0$, $\operatorname{\mathrm{supp}}(\widehat{\varphi})\subset[-1,1]$, $\widehat{\varphi}\geq 0$, $\widehat{\varphi}(0)=1$ and
$\varphi(x)\leq Ce^{-|x|/\sigma(x)}\quad(\text{for all }x\in\mathbb{R})$
for some $C>0$. Let $c>0$ sufficiently small such that $\widehat{\varphi}(x)\geq 1/2$ for $x\in[-c,c]$. The decay condition on $a(\cdot)$ and a routine approximation argument shows the identity
$\int_{\mathbb{R}}\widehat{\varphi}(t-s)\mathrm{d}\mu(t)=\sum_{\lambda\in\mathbb{R}}a(\lambda)\varphi(\lambda)e^{2\pi i\lambda s}$
holds for all $s\in\mathbb{R}$. In particular, there is $C>0$ such that $\mu([s-c,s+c])$ for all $s\in\mathbb{R}$.

### v3

@@ Abstract.
We completely classify Fourier summation formulas, and in particular, all crystalline measures with quadratic decay. Our classification employs techniques from almost periodic functions, Hermite-Biehler functions, de Branges spaces and Poisson representation. We show how our classification generalizes recent results of Kurasov & Sarnak and Olevskii & Ulanovskii. As an application, we give a new classification result for nonnegative measures with uniformly discrete support that are bounded away from zero on their support. Moreover, we give a new construction using eta-quotients, generalizing an old example of Guinand.
@@ 1. Introduction
To understand the atomic structure of a crystal (of some material) one can fire a beam of electrons into the object and observe an electron diffraction pattern on a screen [^1] This technique is called high resolution electron microscopy (HRTEM), but another common technique is X-ray microscopy.. In most materials one observes a periodic diffraction pattern, such as a two-dimensional lattice with $6$-fold symmetry, as in Tantalum pentoxide ${\rm Ta}_{2}{\rm O}_{5}$. However, this is not the case with Holmium-magnesium-zinc Ho-Mg-Zn, where one observes some $5$-fold and $10$-fold symmetries, hence a non-periodic diffraction pattern (see [46, Figures 1 and 2]). The first to record such phenomena experimentally was materials scientist Dan Shechtman, in 1982, and today these materials are called quasicrystals. He was awarded the Nobel Prize in Chemistry in 2011 for his breakthrough. Mathematically, if $\mu$ represents the crystal (so $\mu\geq 0$ and $\operatorname{\mathrm{supp}}(\mu)$ is uniformly discrete) then what we see in the diffraction pattern is effectively $|\widehat{\mu}|^{2}$ (where $\widehat{\mu}$ is the Fourier transform of $\mu$). A famous inverse problem is that of reconstructing atomic structure from diffraction data, overcoming the issue that diffraction only determines intensity of spots but loses phase information. Real-world quasicrystals (such as Decagonite $\text{Al}_{71}\text{Ni}_{24}\text{Fe}_{5}$), although uniformly discrete in physical space, typically have everywhere dense diffraction spectrum [^2] We thank Michael Baake for pointed this out, among several other remarks.. The diffraction pattern is a super-position of Dirac combs with several intensities (Bragg peaks) however, due to instrument precision and tuning, we only see peaks that have, say, an intensity above $10^{-4}$ of the brightest peak. Hence, in real physical experiments, although the diffraction picture visually seems to be locally finite, nevertheless the real measure $\widehat{\mu}$ has a dense support. On the other hand, stable 11D quasicrystals with locally finite spectra do exist in nature, and can be found as projections along pseudo axis’s, see [45, Figure 1] and the references in [19, Ch. 7]. Nowadays there is unified framework to construct measures $\mu$ whose spectrum fits the data observed in the diffraction patterns of physical quasicrystals, these are called model sets (or Meyer sets) [8, 19, 28, 36, 37, 39].
@@
From the mathematical standpoint, many mathematicians have wondered about what different types of measures $\mu$ with pure point support and spectrum are there, and what kind of geometrical properties of the spectrum are allowed under various conditions on $\mu$? The main classification question from the mathematical perspective is:
Can on one classify all measures $\mu$ with pure point support and spectrum?
Freeman Dyson addressed this question in 2009 in his famous paper [18], using the word quasicrystals for such measure, suggesting that such classification might be a very difficult mathematical problem, but made the “outrageous suggestion” it could potentially give a proof of the Riemann Hypothesis [^3] We do not share the same belief.. He said: “My suggestion is the following. Let us pretend that we do not know that the Riemann Hypothesis is true. Let us tackle the problem from the other end. Let us try to obtain a complete enumeration and classification of one-dimensional quasicrystals. That is to say, we enumerate and classify all point distributions that have a discrete point spectrum.” Dyson stated this because of Guinand’s prime summation formula (also known as Weyl’s summation formula), which we explain in the Section 8. The notorious paper of Guinand [22], which inspired Dyson, grasps such classification, but it was forgotten awhile. Meyer resurrected Guinand’s paper in [37].
@@
We say a measure $\mu$ is crystalline if it has locally finite support and its Fourier transform $\widehat{\mu}$ also has locally finite support [^4] Such measures are sometimes also called quasicrystals, Fourier quasicrystals, or doubly sparse, depending on the author, area, context, decay conditions, etc. We will stick with Meyer’s denomination of crystalline measure.. In this scenario, the first geometric classification breakthrough occurred in the work of Lev & Olevskii [30, 31], where they classify classical Poisson summation as the only crystalline measures that have uniformly discrete support and spectrum. Guinand’s example [22, p. 265] shows the theorem is tight (we generalize his construction in Section 8). Recently, the field is very active, and the amazing works of Radchenko & Viazovska [42], Ramos & Sousa [44], Kurasov & Sarnak [26], Olevskii & Ulanovskii [40] and Alon et. al. [2, 1], but also, the not so recent works of Bohr [16], Besicovitch [10], Guinand [22], de Branges [13] and Meyer [35], deeply inspired this paper.
@@
In this paper we completely classify (what we call) Fourier summation formulas, and in particular we classify crystalline measures that have quadratic decay, although our results are more general and can be applied to dense spectra. As far as we know, it is the first time classification results are proven in such generality. Our classification involves heavily the theory of almost periodic analytic functions [10] and de Branges spaces techniques [13]. Almost periodic functions have permeated this are for some time now, for example in the works of Guinand [22], Meyer [38] and Baake et. al. [4, 8, 9], among others. On the other hand, the connection with de Branges spaces that we make here is totally new.
A Fourier summation formula should be one where the identity
@@
holds, at the bear minimum, for all functions $\varphi\in C^{\infty}_{c}(\mathbb{R})$, where $\{\lambda_{n}\}_{n\in\mathbb{N}}\subset\mathbb{R}$ is some sequence of real nodes, $\{a(\lambda_{n})\}_{n\in\mathbb{N}}\subset\mathbb{C}$ are some complex coefficients and $\mu$ is some complex-valued measure. There is a zoo of formulas of such type which we describe in Section 8. We choose the above form of because: (1) We want a summation formula, so there must be a sum somewhere in the definition and such summation must be singular with respect to Lebesgue measure; (2) Allowing the right hand side above to have a vanishing singular part is boring, as otherwise any $L^{1}$-function would give rise to a Fourier summation formula.
@@ 1.1. Overview
Theorems 1 and 2 are our main general classification results and we later specialize them for measures with locally finite support (in particular, also crystalline measures) in Theorems 3 and 4. Finally, we derive a new characterization Theorem 5 for measures with uniformly discrete support such that $\mu|_{\operatorname{\mathrm{supp}}(\mu)}\geq\varepsilon$, but with no geometric assumption on $\widehat{\mu}$, which perhaps makes this result most useful to physics.
@@ 2. Main Results
We will have to introduce several things before we can state our first main result. We say that a complex-valued measure $\mu$ on $\mathbb{R}$ is locally finite if its total variation
@@
is finite on every compact set $K\subset\mathbb{R}$, or equivalently, if there are nonnegative locally finite Borel measures $\mu_{j}$ such that $\mu=\mu_{1}-\mu_{2}+i(\mu_{3}-\mu_{4})$. We say that $\mu$ is tempered if it is a complex-valued locally finite measure and there exists a tempered distribution $u\in\mathcal{S}^{\prime}(\mathbb{R})$ such that $u(\varphi)=\int_{\mathbb{R}}\varphi\mathrm{d}\mu$ for all $\varphi\in C^{\infty}_{c}(\mathbb{R})$. Some authors might not concord with this definition, but we are following the setup of [4]. We say $\mu$ is strongly tempered if $\varphi\in L^{1}(|\mu|)$ for all $\varphi\in\mathcal{S}(\mathbb{R})$. By [4, Prop. 2.5], this is equivalent to the degree of $\mu$
@@
being $\infty$. In particular, the functional $\varphi\mapsto\int_{\mathbb{R}}\varphi\mathrm{d}\mu$ defines a tempered distribution, hence $\mu$ is tempered. The implication [tempered] $\Rightarrow$ [strongly tempered] is false. Indeed, the measure $\mu=\sum_{n\geq 0}3^{n^{2}}({\boldsymbol{\delta}}_{n}-{\boldsymbol{\delta}}_{n+4^{-n^{2}}})$ is not strongly tempered but it is tempered, since $\int 2^{-t^{2}}\mathrm{d}|\mu|(t)\geq\sum_{n\geq 0}3^{n^{2}}2^{-n^{2}}=\infty$ but [^5] Whenever it is convenient, we write $A\ll B$ to mean that there is $C>0$ such that $|A|\leq C|B|$. $|\sum_{n\geq 0}3^{n^{2}}(\varphi({n})-\varphi({n+4^{-n^{2}}}))|\ll\|\varphi^{\prime}\|_{\infty}$.
We say that a function $a:\mathbb{R}\to\mathbb{C}$ is locally summable if $\operatorname{\mathrm{supp}}(a):=\{\lambda\in\mathbb{R}:a(\lambda)\neq 0\}$ is at most countable and for some enumeration $\operatorname{\mathrm{supp}}(a)=\{\lambda_{n}\}_{n\geq 1}$ we have
@@
We say $(\mu,a)$ is a Fourier summation pair $(\mathrm{FS}$-pair)) if $\mu$ is a strongly tempered measure, $a:\mathbb{R}\to\mathbb{C}$ is locally summable and
@@
for $j=1,2$. Since for any $\varphi\in C^{\infty}_{c}(\mathbb{R})$ we have $\varphi=\varphi_{1}-i\varphi_{2}$ where each $\varphi_{j}$ is antipodal, and each $\varphi_{j}={\rm Re}\,\varphi_{j}-i{\rm Im}\,\varphi_{j}$, where ${\rm Re}\,\varphi_{j}$ is real and even and $i{\rm Im}\,\varphi_{j}$ is imaginary and odd (hence both are antipodal as well), we deduce by linearity that $(\mu_{j},a_{j})$ is a real-antipodal $\mathrm{FS}$-pair for $j=1,2$. As an example, note that $\sum_{n\in\mathbb{Z}}i^{n}\widehat{\varphi}(n)=\sum_{n\in\mathbb{Z}}\varphi(n+1/4)$ has the real-antipodal splitting
$\sum_{n\in\mathbb{Z}}(-1)^{n}\widehat{\varphi}(2n)$ $=\frac{1}{2}\sum_{n\in\mathbb{Z}}\left(\varphi(n+1/4)+\varphi(n-1/4)\right)$
$\sum_{n\in\mathbb{Z}}(-1)^{n}\widehat{\varphi}(2n+1)$ $=\frac{1}{2i}\sum_{n\in\mathbb{Z}}\left(\widehat{\varphi}(n+1/4)-\widehat{\varphi}(n-1/4)\right).$
@@
for some $c>0$. We let $\mathbb{C}^{+}=\{z=x+iy:y>0\}$. Following Besicovitch [^6] Besicovitch considers the half-plane ${\rm Re}\,z>0$ instead of ${\rm Im}\,z>0$, so one has to mentally perform a rotation in order to use his results. [10, Ch. 3], we say a holomorphic function $f:\mathbb{C}^{+}\to\mathbb{C}$ is almost periodic if for every $0<\alpha<\beta<\infty$ and $\varepsilon>0$ there is a relatively dense [^7] A set $\tau\subset\mathbb{R}$ is relatively dense if there exists $l>0$ such that $(x,x+l)\cap\tau\neq\emptyset$ for every $x\in\mathbb{R}$. set of $\varepsilon$-translations $\tau\subset\mathbb{R}$ such that
@@
${\rm spec}(f):=\{\lambda\in\mathbb{R}:\mathbb{E}f(\lambda)\neq 0\},$
and it can be shown that this set is at most countable. We say a holomorphic function $g:\mathbb{C}^{+}\to\mathbb{C}$ is of bounded type [^8] This is the same to say that $\log|g(z)|$ has a nonnegative harmonic majorant. if there are two bounded holomorphic functions $P,Q:\mathbb{C}^{+}\to\mathbb{C}$ such that $g=P/Q$. The following is the first main result of this paper.
@@
    $\exp(f)$ is of bounded type.
@@
Note that condition (III) is equivalent to $\mathbb{E}f(\lambda)$ being locally summable only, no growth required.
@@
We claim that if $(\mu,a)$ is an $\mathrm{FS}$-pair such such that $\mu\geq 0$ and $a(\cdot)$ has exponential growth then $\deg(\mu)\leq 2$. Indeed, let $\varphi\in C^{\infty}_{c}(\mathbb{R})$ be even, $0\leq\varphi\leq 1$, $\int\varphi=\varphi(0)=1$ and $\operatorname{\mathrm{supp}}(\varphi)\subset(-1,1)$. Let $\varphi_{\varepsilon}(x)=\varphi(x/\varepsilon)/\varepsilon$ and $\psi_{\varepsilon}(x)=\widehat{\varphi}(x/\varepsilon)/\varepsilon$ for $0<\varepsilon<1$. Let $g(x)=e^{-2\pi C|x|}$, with $C>0$ sufficiently large, and note $\widehat{g}(\xi)=\frac{C}{\pi(C^{2}+\xi^{2})}$. We have
$\int_{\mathbb{R}}(\widehat{g}\widehat{\varphi}_{\varepsilon})*\psi_{\varepsilon}(t-t_{0})\mathrm{d}\mu(t)=\sum_{\lambda\in\mathbb{R}}a(\lambda)g*\varphi_{\varepsilon}(\lambda)\widehat{\psi}_{\varepsilon}(\lambda)e^{-2\pi i\lambda t_{0}},$
for any $t_{0}\in\mathbb{R}$. It is not hard to show that $|g*\varphi_{\varepsilon}(\lambda)\widehat{\psi}_{\varepsilon}(\lambda)|\leq e^{-2\pi C(|\lambda|-1)}$. Thus, we conclude the right hand side above is uniformly bounded in absolute valued for $0<\varepsilon<1$. Since $\mu\geq 0$ and $(\widehat{g}\widehat{\varphi}_{\varepsilon})*\psi_{\varepsilon}\to\widehat{g}$ pointwise, as $\varepsilon\to 0$, we can apply Fatou’s Lemma to deduce $\int\frac{\mathrm{d}\mu(t)}{C^{2}+(t-t_{0})^{2}}<C_{0}$, for some constant $C_{0}>0$ independent of $t_{0}$. In particular we obtain that [^9] Hence $\mu$ is translation-bounded.
$\sup_{t_{0}\in\mathbb{R}}\mu([t_{0}-1,t_{0}+1])\leq(C^{2}+1)C_{0}.$